OpenStudy (anonymous):

This sequence is geometric. What is the next term? 2, 8, 32, 128, … 256 384 428 512

OpenStudy (anonymous):

Same as before :) Look at each pair of terms, and figure out how it gets from the first to the second. It's always the same. Then you just do that to the last number to get the answer. What do you do to 2 to get 8? What do you do to 8 to get 32? What do you do to 32 to get 128? Now you could first try addition: 2+6=8 8+24=32 32+96=128 Ok, it's not addition... what about multiplication? 2X=8 8X=32 32X=128